Janet can paint a kitchen in 3 hours and Charles can paint the same kitchen in 5 hours . How long would it take for both of them working together to paint the kitchen ?
Answer by Alan3354(69443)

Janet can paint a kitchen in 3 hours and Charles can paint the same kitchen in 5 hours . How long would it take for both of them working together to paint the kitchen ?
================
Janet does 1/3 per hour.
Charles does 1/5 per hour.
1/3 + 1/5 = 1/t
-------
3*5/(3+5) = 15/8 hours